OpenStudy (anonymous):

what is the simplest form for 8√63x^5

OpenStudy (anonymous):

is x^5 in root too?

OpenStudy (anonymous):

5 in the exponent of x

OpenStudy (anonymous):

\[8\sqrt{63x^5}\]

OpenStudy (anonymous):

i meant to ask if its in root

OpenStudy (anonymous):

yeah that it

OpenStudy (anonymous):

\[8\sqrt{63x^5} = 8 \sqrt{9*7x^4*x} = 8*3x^2\sqrt{7x} = 24x^2\sqrt{7x}\]

OpenStudy (anonymous):

so \[24^{2}\sqrt{7}x\]

OpenStudy (anonymous):

is the answer ?

OpenStudy (anonymous):

i meant to put an x after 24

OpenStudy (anonymous):

24 x^3 * and yes thats it

OpenStudy (anonymous):

Thanks <3

OpenStudy (anonymous):

you are welcome :)